### v2

@@ 8. $J$-invariant for groups of outer type
where ``"``"`$\ldots$"` ` … " stands for a sum of monomials of degree $\geq 3$. Therefore, we have $P(\mathop{\mathrm{CH}}\nolimits^{*}_{K}(U),t)\cdot P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(X),t)=1+t+2t^{2}+\ldots\,,$ and this contradicts Proposition 7.7. Thus, $j_{2}=0$.
@@
Recall that $[K:F]=p^{n}$ (where $K$ is the field from the definition of normed Chow rings such that the group $G_{K}$ is of inner type) for some prime number $p$ and for some $n\geq 1$. If the group $G$ is split over $L$, then by the properties of the Weil restrictions (see Section 3) the functor $R_{L/F}$ induces an isomorphism $\mathop{\mathrm{Ch}}\nolimits^{*}(G/B)\xrightarrow{\simeq}\mathop{\mathrm{CH}}\nolimits^{*}_{K}(R_{L/F}(G/B))$. Finally, if $G$ is a not split over $L$, then the functor $R_{L/F}$ induces an isomorphism $\mathop{\mathrm{CH}}\nolimits^{*}_{K}(G/B)\xrightarrow{\simeq}\mathop{\mathrm{CH}}\nolimits^{*}_{K}(R_{L/F}(G/B))$.
@@
${}^{2}\mathrm{A}_{2n+1}$ $2$ $n+2$ $12,\,i=$ 1 $2^{k_{1}}\parallel(n+1)$ the same restrictions
$2i-3,\,i\geq 2$ $1$ assuming $i,m\geq 2$
${}^{2}\mathrm{D}_{n}$ $2$ $[\frac{n+1}{2}]+1$ $11,\,i=$ 1 $\tfrac{1+(-1)^{n+1}}{2}$ if $d_{i}+l=2^{s}d_{m}$ and
$22,\,i=$ 2 $[\log_{2}n]$ $2\nmid\binom{d_{i}}{l}$, then $j_{m}\leq j_{i}+s$
$2i-3,\,i\geq 3$ $[\log_{2}\frac{2n}{2i-3}]$ assuming $i,m\geq 2$
${}^{3}\mathrm{D}_{4}$ $3$ $1$ $4$ $1$
${}^{2}\mathrm{E}_{6}$ $2$ $3$ 3, 5, 93593,\,5,\,93 , 5 , 9 1, 1, $,\,1,\,1$ 1 $j_{1}\geq j_{2}\geq j_{3}$
$R_{L/F}(G)$ $p$ $r(G)$ $[L:F]\cdot d_{i}(G)$ $k_{i}(G)$ as for $G$
@@
Let $G_{0}$ be a quasi-split semisimple algebraic group over a field $F$, let $K$ be a minimal field extension such that the group $(G_{0})_{K}$ is of inner type, let $B$ be a Borel subgroup of $G_{0}$, let $E$ be a $G_{0}$-torsor over $F$ and let $p$ be a torsion prime of $G_{0}$. Assume that $[K:F]=p^{n}$ for some $n\geq 1$.

### v3

@@ 8. $J$-invariant for groups of outer type
where $\ldots$" stands for a sum of monomials of degree $\geq 3$. Therefore, we have
$P(\mathop{\mathrm{CH}}\nolimits^{*}_{K}(U),t)\cdot P(\overline{\mathop{\mathrm{CH}}\nolimits}_{K}^{*}(X),t)=1+t+2t^{2}+\ldots\,$,
and this contradicts Proposition 7.7. Thus, $j_{2}=0$.
@@
Recall that $K:F]=p^{n}$ (where $K$ is the field from the definition of normed Chow rings such that the group $G_{K}$ is of inner type) for some prime number $p$ and for some $n\geq 1$. If the group $G$ is split over $L$, then by the properties of the Weil restrictions (see Section 3) the functor $R_{L/F}$ induces an isomorphism $\mathop{\mathrm{Ch}}\nolimits^{*}(G/B)\xrightarrow{\simeq}\mathop{\mathrm{CH}}\nolimits^{*}_{K}(R_{L/F}(G/B))$. Finally, if $G$ is a not split over $L$, then the functor $R_{L/F}$ induces an isomorphism $\mathop{\mathrm{CH}}\nolimits^{*}_{K}(G/B)\xrightarrow{\simeq}\mathop{\mathrm{CH}}\nolimits^{*}_{K}(R_{L/F}(G/B))$.
@@
$^{2}\mathrm{A}_{2n+1}$ 22 $n+2$ 2,i=12,\,i=1 $2^{k_{1}}\parallel(n+1)$ the same restrictions
$2i-3,\,i\geq 2$ 11 assuming $i,m\geq 2$
$^{2}\mathrm{D}_{n}$ 22 $\frac{n+1}{2}]+1$ 1,i=11,\,i=1 $\tfrac{1+(-1)^{n+1}}{2}$ if $d_{i}+l=2^{s}d_{m}$ and
2,i=22,\,i=2 $\log_{2}n]$ $2\nmid\binom{d_{i}}{l}$, then $j_{m}\leq j_{i}+s$
$2i-3,\,i\geq 3$ $\log_{2}\frac{2n}{2i-3}]$ assuming $i,m\geq 2$
$^{3}\mathrm{D}_{4}$ 33 11 44 11
$^{2}\mathrm{E}_{6}$ 22 33 3, 5, 93,\,5,\,9 1, 1, 11,\,1,\,1 $j_{1}\geq j_{2}\geq j_{3}$
$R_{L/F}(G)$ $p$ $r(G)$ $L:F]\cdot d_{i}(G)$ $k_{i}(G)$ as for $G$
@@
Let $G_{0}$ be a quasi-split semisimple algebraic group over a field $F$, let $K$ be a minimal field extension such that the group $(G_{0})_{K}$ is of inner type, let $B$ be a Borel subgroup of $G_{0}$, let $E$ be a $G_{0}$-torsor over $F$ and let $p$ be a torsion prime of $G_{0}$. Assume that $K:F]=p^{n}$ for some $n\geq 1$.